two mens with their weights in the ratio 5:3 run up a staircase in times in the ratio 11:9.Then the ratio of power of first to that second is

Grade:10

2 Answers

Arun
6 years ago
As we know
Power (P) = m*g*h/t
or
\alpha m/t
Hence
P1/P2 = m1* t2 /m2 *t1
Hence
P1/P2 = 5*9/3*11 = 45/33 = 15/11
 
Hope it helps
Nikhil Agarwal
5 years ago
Below is the answerPower = work/tHere the work= mgh Therefore power = mgh/tHere g and h will be same for both men let these are x and y respectively Now P1/P2= m1xy/t1 ÷ m2xy/t2P1/P2= m1t2/m2t1P1/P2= 5×9/3×11P1/P2=45/33P1/P2=15/11= 15:11